give solutions for (x-7)^2=18 give solutions not the decimals
Answer by stanbon(75887)

(x-7)^2=18
------
(x-7) = sqrt(18) or x-7 = -sqrt(18)
--------
x = 7 + 3sqrt(2) or x = 7 - 3sqrt(2)
-----------------------------------------
Cheers,
Stan H.
============